OpenStudy (anonymous):

Solve this equation on the interval [0,2π) 2cos(x/2)-√3=0 Help please & Thank you!!!

OpenStudy (anonymous):

start with \[\cos(\frac{x}{2})=\frac{\sqrt3}{2}\] then solve for \(\frac{x}{2}\) then double to solve for \(x\)

OpenStudy (anonymous):

ok. gotcha! thank you so much! i appreciate it!

OpenStudy (anonymous):

am i supposed to end up with : cos=2√3/2x ???

OpenStudy (anonymous):

heck no

OpenStudy (anonymous):

my bad...

OpenStudy (anonymous):

where do i go from where you told me?

OpenStudy (anonymous):

you are supposed to get \(\frac{x}{2}=\frac{\pi}{6}\)

OpenStudy (anonymous):

since \[\cos(\frac{\pi}{6})=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

hold up, where did the pi come from?

OpenStudy (anonymous):

you working in degrees or radians?

OpenStudy (anonymous):

the question does not specify..

OpenStudy (anonymous):

if you are working in degrees, then since \[\cos(30)=\frac{\sqrt3}{2}\]you get \[\frac{x}{2}=30\iff x=60\]

OpenStudy (anonymous):

ohhhh, i THINK i got it! thank you so much!